# Short remarks on shallow unitary circuits

### v4

@@ 1 Symmetric shallow circuit $\neq$ symmetric global random unitary
Recently it has been shown [16, 21] that an ensemble of random local quantum circuits $U$ on a one-dimensional chain of $n$ qudits of depth $k\operatorname{polylog}(nk)$ is indistinguishable from the Haar random global unitary on $n$ qubits if we query $U$ no more than $k$ times. [^1] The indistinguishability can also be established under a cryptographic assumption with a milder restriction on the number of queries to the unitary [21]. Such an ensemble of unitaries is called a unitary $k$-design. The existence of a shallow unitary design is rather surprising since the depth is much smaller than the linear system size $n$, so the absolute lightcone of an observable at one end of the chain does not even reach the other end. Nonetheless, the ensemble contains sufficient randomness which prohibits any predetermined correlation function from probing the difference between Haar random global unitaries and shallow local circuits.
@@
where $Q_{i}$ is a nonzero hermitian operator supported on a qudit $i$. We assume that every $Q_{i}$ has an eigenvalue zero. [^3] One may further require that each $Q_{i}$ have an integer eigenvalue spectrum so that $e^{{\mathbf{i}}\theta}\mapsto e^{{\mathbf{i}}\theta Q}$ is a representation of $\mathsf{U}(1)$}, but this condition is not used in our argument. This conditions is trivially fulfilled by adding a real multiple of identity to $Q_{i}$. Let ${\mathsf{S}}_{i,j}={\mathsf{S}}_{j,i}$ be the swap operator that interchanges qudit $i$ and $j$. For simplicity we assume that $Q_{i}={\mathsf{S}}_{i,j}Q_{j}{\mathsf{S}}_{i,j}^{\dagger}$. The symmetric unitary group of interest consists of all $U\in{\mathsf{U}}(p^{n})$ that commutes with $Q$. A symmetric local unitary circuit is a composition of layers of nonoverlapping nearest-neighbor 2-qudit unitaries, each of which commutes with $Q$. Of course, any symmetric local unitary circuit belongs to the symmetric unitary group. [^4] The converse is false; there exists a symmetric unitary that is not expressible by any symmetric local unitary circuit [17]. This is however unimportant since we focus on a lower bound on the depth, which is still meaningful since symmetric local unitary circuits with sufficient depths can form a symmetric unitary $k$-design for $k$ that grows with the system size [11]. The depth is the number of layers. This is a standard setting in the context of symmetry-protected topological phases [3].
@@
Since our protocol uses predetermined measurements, it is at least as strong a notion of distinguishability as the additive error for unitary designs, which is more of an information-theoretic measure. Under the latter notion where two mixed unitary channels are sufficiently far apart in the diamond distance, input states and output measurements for the task of distinguishing channels may be implicit and difficult to find.
This completes the analysis of the protocol. In conclusion, for any ensemble of symmetric local circuits to form an approximate symmetric unitary $k$-design with $k\geq 1$ and a small constant additive error, the depth of the circuit must be at least linear in the linear system size. A few comments are below.
The randomization by $U_{AB}$ is important to make the protocol sound. If we omitted it, the measurement of $Q_{A}$ may give a misleading answer for the distinguishing task because symmetric circuits can conspire to transport an abelian charge from anywhere to anywhere instantaneously by creating and annihilating charge-anticharge pairs.
@@
The existence of a conserved total charge $Q$ is implied by a continuous symmetry, but that is not the only source. The conserved quantity $Q$ can be a local Hamiltonian itself, in which case our protocol is simply measuring the energy density. [^5] Some other anticoncentration measures [23] were numerically observed to grow qualitatively differently depending on whether the dynamics conserves energy. Even with a discrete symmetry, a conserved quantity that is a sum of local operators may appear in the following “dilute symmetry” scenario. Suppose that we have a comb [^6] I thank Ehud Altman for suggesting a comb to explain this dilute symmetry. connectivity graph as shown in Figure 2. A qudit is placed on every vertex, and 2-qudit local unitaries are allowed only over the edges. Label each qudit by $(i,\text{tip})$ or $(i,\text{spine})$. We impose a symmetry by a tensor product operator $Q^{\prime}=\prod_{i}X_{(i,\text{tip})}$. Here the symbol $X$ is reminiscent of Pauli $X$ as in the onsite ${\mathbb{Z}}_{2}$ symmetry on qubits. Then, every local unitary gate sees at most one tensor factor of $Q^{\prime}$ and hence $Q=\sum_{i}X_{(i,\text{tip})}+X_{(i,\text{tip})}^{\dagger}$, in addition to $Q^{\prime}$, is preserved by any symmetric local circuit dynamics. A symmetric global unitary is assumed to preserve $Q^{\prime}$ but not necessarily $Q$. This relaxed symmetry condition only widens the gap in the expectation values of $Q_{A}$ and the protocol above still disambiguates shallow symmetric local circuits from global symmetric Haar random unitaries.
We have assumed that the sizes of subsystems $|A|$ and $|B|$ are equal, that the total system is a one-dimensional array of qudits, and that all the local operators $Q_{i}$ are the same up to relabeling of the qudits. These assumptions are for simplicity in presentation and are unimportant. The subsystem $A$ is where we put some known value of charge. If $A$ and $B$ are comparable in size, then a symmetric Haar random unitary $U_{AB}$ on $AB$ will evenly distribute the charge over $AB$ and $A$ will retain a significant fraction of the initial charge. In the symmetric shallow circuit case, the measurement of $Q_{A}$ will reveal this large fraction of the inital charge, but in the global symmetric Haar random case, the measured value of $Q_{A}$ will be suppressed by a factor of the overall system size. It remains to investigate the case of discrete symmetries represented by product operators that are not necessarily dilute.
@@ 2 Shallow circuit learning
For the last step, we recall a simple strategy that appeared in [12] where it was necessary to estimate $\operatorname{{Tr}}(UXU^{\dagger}Y)$ for various local (Pauli) operators $X$ and $Y$ given query access to $U$. Take any density matrix $\rho\succeq 0$ that, as an operator, is supported on some small set of qudits. This mixed state $\rho$ can be prepared by some local quantum channel on its support and some other channel that prepares all the qudits elsewhere in a random product state. The conjugate $U\rho U^{\dagger}$ is a density operator supported on a slightly bigger set of qudits, and usual state tomography can determine $U\rho U^{\dagger}$ whose complexity does not depend on the system size but only on the depth of UU. [^7] If $U$ is a shallow circuit consisting of gates from a discrete set (e.g., Clifford ++ T), then the complexity depends only on the depth. If the gates comprising $U$ is from, say, ${\mathsf{SU}}(4)$, then $U\rho U^{\dagger}$ can only be determined up to some error with some high confidence, where the accuracy and confidence must be boosted for various global requirements, and system size dependence may become necessary. If the tomography has $\sim\epsilon^{-2}\log(1/\delta)$ sample complexity for accuracy $\epsilon$ in some state distance and confidence $1-\delta$, then one would want to consume $\sim n^{2}\epsilon^{-2}\log(n/\delta)$ samples of $U\rho U^{\dagger}$ to ensure that the global accuracy of $U\otimes U^{\dagger}$ is at most $\epsilon$ with confidence $1-\delta$. We can use this to learn $(U\otimes{\mathbf{1}}){\mathsf{S}}_{i,i^{\prime}}(U^{\dagger}\otimes{\mathbf{1}})$ even though the unitary ${\mathsf{S}}_{i,i^{\prime}}$ is not a density operator. Any operator $O=\frac{O+O^{\dagger}}{2}+{\mathbf{i}}\frac{O-O^{\dagger}}{2{\mathbf{i}}}$ is a linear combination of two hermitian operators, and any hermitian operator $H$ is a linear combination of two normalized density operators:
@@
where $\alpha=\sum_{j:\lambda_{j}>0}\lambda_{j}$ and $\beta=\sum_{j:\lambda_{j}<0}|\lambda_{j}|$. Hence, $O$ is a linear combination of four density operators. Since this linear combination is known beforehand, we determine $(U\otimes{\mathbf{1}}){\mathsf{S}}_{i,i^{\prime}}(U^{\dagger}\otimes{\mathbf{1}})$ to any desired accuracy (in operator norm) by local preparations and measurements and oracular access to $U$. No $U^{\dagger}$ is needed. In fact, it is unnecessary to consider the doubled system for this step. It suffices to determine $UX_{i}U^{\dagger}$ (to accuracy $O(\epsilon/n)$ in operator norm for system of $n$ sites) for single-qudit operators $X_{i}$ that generate the single-qudit operator algebra. Since a matrix algebra is generated by just two operators (e.g. $|\sum_{j\in{\mathbb{Z}}/p}\ket{j+1}\!\bra{j}$ and $|\sum_{j\in{\mathbb{Z}}/p}e^{2\pi{\mathbf{i}}j/p}\ket{j}\!\bra{j}$), this means that we have to run a local state tomography for no more than eight states. If each site is a qubit, it suffices to determine just two density operators $U\frac{{\mathbf{1}}+\sigma^{x}_{i}}{2}U^{\dagger}=\frac{1}{2}U\sigma^{x}_{i}U^{\dagger}+\frac{1}{2}{\mathbf{1}}$ and $U\frac{{\mathbf{1}}+\sigma^{z}_{i}}{2}U^{\dagger}$ for each $i$, which are $U$-evolved Pauli stabilizer states. The qubit case is also found in [13, 24].
@@
QCA.    We used only the condition that $U$ maps a local operator to a local operator in the above algorithm, and hence $U$ could have been any quantum cellular automaton (QCA) — by definition, a QCA $\alpha$ is an automorphism [^8] If the operator algebra is finite dimensional, then the Skolem–Noether theorem says that there exists a unitary $U$ such that $\alpha(X)=UXU^{\dagger}$ for all operators $X$. Hence, it is safe to regard $\alpha$ as some global unitary. of an operator algebra such that there exists a constant $r>0$ such that for any single-site operator $X$, the image $\alpha(X)$ is supported on the $r$-neighborhood of the support of $X$. The minimal constant $r$ here is the spread of a QCA. If a circuit has depth $d$, then we know that its spread (of the QCA defined by the conjugation by the circuit) is $r=O(d)$. A primary question in the context of QCA is the converse: given an automorphism $\alpha$ of a local operator algebra of spread $r$, does there exist a quantum circuit of depth $d$ that implements $\alpha$? If so, how does $d$ depend on the system size and the spread? If $d$ is a constant independent of the system size, then the QCA is said to be trivial. It is quite a widely open problem how to determine whether a given QCA, or an infinite sequence of QCA of a fixed spread on finite systems of diverging sizes, is trivial. [^9] This is well understood in one dimension with qudits without tails [7], with tails [20], with fermions [6], or with symmetry [18]; and in two dimensions with qudits without tails [4] (see also [8]); and in any dimensions in the translation invariant prime Clifford category [10], but other situations remain open. See the introduction of [10] and references therein. The above algorithm outputs a shallow circuit that implements $\alpha\otimes\alpha^{-1}$, but does not address the QCA triviality problem. A related question for the circuit learning problem is if it is possible to output a circuit (of a similar depth to the input circuit’s) that does not use any ancillas. Interestingly, a related state learning problem uses an arbitrarily small density of ancillas to output a circuit that creates a given shallow-circuit-evolved product state [15].
@@ 3 Gate complexity of QCA
Here, we present another proof of the theorem of [2] by invertible subalgebra pumping. Our proof is inductive in the spatial dimension $D$ where the base case is D=1D=1.111111We could take $D=0$ as the base case, but it would be more instructive to start with $D=1$.
@@
Given an invertible subalgebra $\mathcal{A}$} of $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D-1})$, we construct an associated QCA in $D$ dimensions as follws. Consider an (infinite) stack of (D1)(D-1)-dimensional sheets. Each sheet supports a copy of $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D-1})$. The full algebra of the stack, $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D})$, is the (infinite) tensor product of those of the sheets. To define an algebra endomorphism on $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D})$, it suffices to specify the images of generators, which we take to be single-qudit operators. A single-qudit operator $x$ is supported on, say, the $k$-th sheet, and we have a decomposition $x^{(k)}=\sum_{j}a_{j}^{(k)}b_{j}^{(k)}$ according to the definition of $\mathcal{A}$}. Declare that $x^{(k)}\mapsto\sum_{j}a_{j}^{(k)}b_{j}^{(k+1)}$, where $b_{j}$ are sent to the next sheet. It can be verified [9, 3.9] that this is indeed an algebra automorphism that preserves the locality with a spread bounded by a function of $r$ that appears in the definition of $\mathcal{A}$}. We denote by $\beta_{\mathcal{A}}$ the QCA defined from $\mathcal{A}$}.
@@
In connection to the previous section on the shallow circuit learning problem, an open problem is whether there is an efficient query algorithm for the conversion from a QCA to a staircase circuit. The two types of data that define a QCA $\alpha$ — one a table of images of single-qudit operators under a QCA and the other a local unitary circuit in a staircase fasion — contain the same number $O(n)$ of parameters up to a fixed multiplicative constant. The former is perhaps the most obvious type of data for $\alpha$, but the latter is more appropriate if we are to apply $\alpha$ to some arbitrary operator whose decomposition in terms of single-qudit operators is complicated.

### v5

@@ 1 Symmetric shallow circuit $\neq$ symmetric global random unitary
Recently it has been shown [17, 22] that an ensemble of random local quantum circuits $U$ on a one-dimensional chain of $n$ qubits of depth $k\operatorname{polylog}(nk)$ is indistinguishable from the Haar random global unitary on $n$ qubits if we query $U$ no more than $k$ times. [^1] The indistinguishability can also be established under a cryptographic assumption with a milder restriction on the number of queries to the unitary [22]. Such an ensemble of unitaries is called a unitary $k$-design. The existence of a shallow unitary design is rather surprising since the depth is much smaller than the linear system size $n$, so the absolute lightcone of an observable at one end of the chain does not even reach the other end. Nonetheless, the ensemble contains sufficient randomness which prohibits any predetermined correlation function from probing the difference between Haar random global unitaries and shallow local circuits.
@@
where $Q_{i}$ is a nonzero hermitian operator supported on a qudit $i$. We assume that every $Q_{i}$ has an eigenvalue zero. [^3] One may further require that each $Q_{i}$ have an integer eigenvalue spectrum so that $e^{{\mathbf{i}}\theta}\mapsto e^{{\mathbf{i}}\theta Q}$ is a representation of $\mathsf{U}(1)$}, but this condition is not used in our argument. This condition is trivially fulfilled by adding a real multiple of identity to $Q_{i}$. Let ${\mathsf{S}}_{i,j}={\mathsf{S}}_{j,i}$ be the swap operator that interchanges qudit $i$ and $j$. For simplicity we assume that $Q_{i}={\mathsf{S}}_{i,j}Q_{j}{\mathsf{S}}_{i,j}^{\dagger}$. The symmetric unitary group of interest consists of all $U\in{\mathsf{U}}(p^{n})$ that commutes with $Q$. A symmetric local unitary circuit is a composition of layers of nonoverlapping nearest-neighbor 2-qudit unitaries, each of which commutes with $Q$. Of course, any symmetric local unitary circuit belongs to the symmetric unitary group. [^4] The converse is false; there exists a symmetric unitary that is not expressible by any symmetric local unitary circuit [18]. This is however unimportant since we focus on a lower bound on the depth, which is still meaningful since symmetric local unitary circuits with sufficient depths can form a symmetric unitary $k$-design for $k$ that grows with the system size [11]. The depth is the number of layers. This is a standard setting in the context of symmetry-protected topological phases [3].
@@
Since our protocol is explicit, it is at least as strong a notion of distinguishability as the additive error for unitary designs, which is more of an information-theoretic measure. Under the latter notion where two mixed unitary channels are sufficiently far apart in the diamond distance, input states and output measurements for the task of distinguishing channels may be implicit and difficult to find.
This completes the analysis of the protocol. In conclusion, for any ensemble of symmetric local circuits to form an approximate symmetric unitary $k$-design with $k\geq 1$ and a small constant additive error, the depth of the circuit must be at least linear in the linear system size. Some comments are below.
The randomization by $U_{AB}$ is important to make the protocol sound. If we omitted it, the measurement of $Q_{A}$ may give a misleading answer for the distinguishing task because symmetric circuits can conspire to transport an abelian charge from anywhere to anywhere instantaneously by creating and annihilating charge-anticharge pairs. [^5] Note that $U_{AB}$ is the only component for which the implementation gate complexity as a function of $n$ might be high; other components beyond querying a blackbox $W$ are preparation of eigenstates of $Q_{i}$ and measurement of $Q_{i}$, both of which are local operations. Since the protocol is to measure $\overline{Q_{A}}=\operatorname{{\mathbb{E}}}_{U_{AB},W}\operatorname{{Tr}}(Q_{A}WU_{AB}\rho_{0}U_{AB}^{\dagger}W^{\dagger})$ with a small additive error, we only need $U_{AB}$ to be drawn from an approximate symmetric unitary $1$-design. It appears that the lower bound on the convergence rate of $U(1)$-symmetric circuits to $U(1)$-symmetric approximate designs is largely an open problem.
@@
The existence of a conserved total charge $Q$ is implied by a continuous symmetry, but that is not the only source. The conserved quantity $Q$ can be a local Hamiltonian itself, in which case our protocol is simply to measure the energy density. [^6] Some other anticoncentration measures [24] were numerically observed to grow qualitatively differently depending on whether the dynamics conserves energy. Even with a discrete symmetry, a conserved quantity that is a sum of local operators may appear in the following “dilute symmetry” scenario. Suppose that we have a comb [^7] I thank Ehud Altman for suggesting a comb to explain this dilute symmetry. connectivity graph as shown in Figure 2. A qudit is placed on every vertex, and 2-qudit local unitaries are allowed only over the edges. Label each qudit by $(i,\text{tip})$ or $(i,\text{spine})$. We impose a symmetry by a tensor product operator $Q^{\prime}=\prod_{i}X_{(i,\text{tip})}$. Here the symbol $X$ is reminiscent of Pauli $X$ as in the onsite ${\mathbb{Z}}_{2}$ symmetry on qubits. Then, every local unitary gate sees at most one tensor factor of $Q^{\prime}$ and hence $Q=\sum_{i}X_{(i,\text{tip})}+X_{(i,\text{tip})}^{\dagger}$, in addition to $Q^{\prime}$, is preserved by any symmetric local circuit dynamics. A symmetric global unitary is assumed to preserve $Q^{\prime}$ but not necessarily $Q$. This relaxed symmetry condition only widens the gap in the expectation values of $Q_{A}$ and the protocol above still disambiguates shallow symmetric local circuits from global symmetric Haar random unitaries.
We have assumed that the sizes of subsystems $|A|$ and $|B|$ are equal, that the total system is a one-dimensional array of qudits, and that all the local operators $Q_{i}$ are the same up to relabeling of the qudits. These assumptions are for simplicity in presentation and are unimportant. The subsystem $A$ is where we put some known value of charge. If $A$ and $B$ are comparable in size, then a symmetric Haar random unitary $U_{AB}$ on $AB$ will evenly distribute the charge over $AB$ and $A$ will retain a significant fraction of the initial charge. In the symmetric shallow circuit case, the measurement of $Q_{A}$ will reveal this large fraction of the initial charge, but in the global symmetric Haar random case, the measured value of $Q_{A}$ will be suppressed by a factor of the overall system size. It remains to investigate the case of discrete symmetries represented by product operators that are not necessarily dilute.
@@ 2 Shallow circuit learning
For the last step, we recall a simple strategy that appeared in [12] where it was necessary to estimate $\operatorname{{Tr}}(UXU^{\dagger}Y)$ for various local (Pauli) operators $X$ and $Y$ given query access to $U$. Take any density matrix $\rho\succeq 0$ that, as an operator, is supported on some small set of qudits. This mixed state $\rho$ can be prepared by some local quantum channel on its support and some other channel that prepares all the qudits elsewhere in a random product state. The conjugate $U\rho U^{\dagger}$ is a density operator supported on a slightly bigger set of qudits, and usual state tomography can determine $U\rho U^{\dagger}$ whose complexity does not depend on the system size but only on the depth of UU. [^8] If $U$ is a shallow circuit consisting of gates from a discrete set (e.g., Clifford ++ T), then the complexity depends only on the depth. If the gates comprising $U$ are from, say, ${\mathsf{SU}}(4)$, then $U\rho U^{\dagger}$ can only be determined up to some error with some high confidence, where the accuracy and confidence must be boosted for various global requirements, and system size dependence may become necessary. If the tomography has $\sim\epsilon^{-2}\log(1/\delta)$ sample complexity for accuracy $\epsilon$ in some state distance and confidence $1-\delta$, then one would want to consume $\sim n^{2}\epsilon^{-2}\log(n/\delta)$ samples of $U\rho U^{\dagger}$ to ensure that the global accuracy of $U\otimes U^{\dagger}$ is at most $\epsilon$ with confidence $1-\delta$. We can use this to learn $(U\otimes{\mathbf{1}}){\mathsf{S}}_{i,i^{\prime}}(U^{\dagger}\otimes{\mathbf{1}})$ even though the unitary ${\mathsf{S}}_{i,i^{\prime}}$ is not a density operator. Write ${\mathsf{S}}_{i,i^{\prime}}=\frac{1}{p}\sum_{P}P_{i}\otimes P^{\dagger}_{i^{\prime}}$ as a linear combination of (generalized) Pauli operators, where $p$ is the local qudit dimension. Clearly, it suffices to learn $UPU^{\dagger}$ for every (generalized) Pauli operator $P$, but since $P\mapsto UPU^{\dagger}$ is an algebra homomorphism, it suffices to learn just two operators $UXU^{\dagger}$ and $UZU^{\dagger}$ where $|X=\sum_{j\in{\mathbb{Z}}/p}\ket{j+1}\!\bra{j}$ and $|Z=\sum_{j\in{\mathbb{Z}}/p}e^{2\pi{\mathbf{i}}j/p}\ket{j}\!\bra{j}$. Any operator $O=\frac{O+O^{\dagger}}{2}+{\mathbf{i}}\frac{O-O^{\dagger}}{2{\mathbf{i}}}$ is a linear combination of two hermitian operators, and any hermitian operator $H$ is a linear combination of two normalized density operators:
@@
where $\alpha=\sum_{j:\lambda_{j}>0}\lambda_{j}$ and $\beta=\sum_{j:\lambda_{j}<0}|\lambda_{j}|$. Hence, $O$ is a linear combination of four density operators. It follows that, by multiplying and summing the results of tomography of eight local states, we infer $(U\otimes{\mathbf{1}}){\mathsf{S}}_{i,i^{\prime}}(U^{\dagger}\otimes{\mathbf{1}})$. The needed accuracy of the tomography is $O(\epsilon/n\operatorname{poly}(p))$ in operator norm to ensure that the error in the reconstructed $U\otimes U^{\dagger}$ is at most $\epsilon$. If each site is a qubit, it suffices to determine just two density operators $U\frac{{\mathbf{1}}+\sigma^{x}_{i}}{2}U^{\dagger}=\frac{1}{2}U\sigma^{x}_{i}U^{\dagger}+\frac{1}{2}{\mathbf{1}}$ and $U\frac{{\mathbf{1}}+\sigma^{z}_{i}}{2}U^{\dagger}$ for each $i$, which are $U$-evolved Pauli stabilizer states. The qubit case is also found in [13, 25].
@@
QCA.    We used only the condition that $U$ maps a local operator to a local operator in the above algorithm, and hence $U$ could have been any quantum cellular automaton (QCA) — by definition, a QCA $\alpha$ is an automorphism [^9] If the operator algebra is finite dimensional, then the Skolem–Noether theorem says that there exists a unitary $U$ such that $\alpha(X)=UXU^{\dagger}$ for all operators $X$. Hence, it is safe to regard $\alpha$ as some global unitary. of an operator algebra such that there exists a constant $r>0$ such that for any single-site operator $X$, the image $\alpha(X)$ is supported on the $r$-neighborhood of the support of $X$. The minimal constant $r$ here is the spread of a QCA. If a circuit has depth $d$, then we know that its spread (of the QCA defined by the conjugation by the circuit) is $r=O(d)$. A primary question in the context of QCA is the converse: given an automorphism $\alpha$ of a local operator algebra of spread $r$, does there exist a quantum circuit of depth $d$ that implements $\alpha$? If so, how does $d$ depend on the system size and the spread? If $d$ is a constant independent of the system size, then the QCA is said to be trivial. It is quite a widely open problem how to determine whether a given QCA, or an infinite sequence of QCA of a fixed spread on finite systems of diverging sizes, is trivial. [^10] This is well understood in one dimension with qudits without tails [7], with tails [21], with fermions [6], or with symmetry [19]; and in two dimensions with qudits without tails [4] (see also [8]); and in any dimensions in the translation invariant prime Clifford category [10], but other situations remain open. See the introduction of [10] and references therein. The above algorithm outputs a shallow circuit that implements $\alpha\otimes\alpha^{-1}$, but does not address the QCA triviality problem. A related question for the circuit learning problem is if it is possible to output a circuit (of a similar depth to the input circuit’s) that does not use any ancillas. Interestingly, a related state learning problem uses an arbitrarily small density of ancillas to output a circuit that creates a given shallow-circuit-evolved product state [16, 14].
@@ 3 Gate complexity of QCA
Here, we present another proof of the theorem of [2] by invertible subalgebra pumping. Our proof is inductive in the spatial dimension $D$ where the base case is D=1D=1.121212We could take $D=0$ as the base case, but it would be more instructive to start with $D=1$.
@@
Given an invertible subalgebra $\mathcal{A}$} of $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D-1})$, we construct an associated QCA in $D$ dimensions as follows. Consider an (infinite) stack of (D1)(D-1)-dimensional sheets. Each sheet supports a copy of $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D-1})$. The full algebra of the stack, $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D})$, is the (infinite) tensor product of those of the sheets. To define an algebra endomorphism on $\operatorname*{{\mathsf{Mat}}}(\Lambda_{D})$, it suffices to specify the images of generators, which we take to be single-qudit operators. A single-qudit operator $x$ is supported on, say, the $k$-th sheet, and we have a decomposition $x^{(k)}=\sum_{j}a_{j}^{(k)}b_{j}^{(k)}$ according to the definition of $\mathcal{A}$}. Declare that $x^{(k)}\mapsto\sum_{j}a_{j}^{(k)}b_{j}^{(k+1)}$, where $b_{j}$ are sent to the next sheet. It can be verified [9, 3.9] that this is indeed an algebra automorphism that preserves the locality with a spread bounded by a function of $r$ that appears in the definition of $\mathcal{A}$}. We denote by $\beta_{\mathcal{A}}$ the QCA defined from $\mathcal{A}$}.
@@
In connection to the previous section on the shallow circuit learning problem, an open problem is whether there is an efficient query algorithm for the conversion from a QCA to a staircase circuit. The two types of data that define a QCA $\alpha$ — one a table of images of single-qudit operators under a QCA and the other a local unitary circuit in a staircase fashion — contain the same number $O(n)$ of parameters up to a fixed multiplicative constant. The former is perhaps the most obvious type of data for $\alpha$, but the latter is more appropriate if we are to apply $\alpha$ to some arbitrary operator whose decomposition in terms of single-qudit operators is complicated.